# Graph-Structured Topic Modeling for Documents with Spatial or Covariate Dependencies

### v1

@@ C.3.2 Technical Lemmas
$\begin{split}&\mathbb{P}\left[\exists k\in[K],\exists l\in[n_{C}]:\quad\frac{1}{n_{\mathcal{C}_{l}}N}|\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{N}\sum_{j=1}^{p}\eta_{jm}|>\sqrt{C^{*}\frac{\log(n)}{n_{\mathcal{C}_{l}}N}}\right]\\ &\leq 2Kn_{C}\exp\{-{C^{*}\log(n)}\}\\ &=\exp\{\log(2)+\log(K)+\log(n_{C})-C^{*}\log(n)\}\\ &\leq\exp\{-(C-3)\log(n)\},\end{split}$
where the last line holds for some large enough constant $\tilde{C}<C^{*}$ such that $n^{\tilde{C}}\geq 2Kn_{C}$ and $C=C^{*}-\tilde{C}$. Thus with probability $1-o(n^{-1})$, for all $k\in[K]$ and all $l\in[n_{C}]$:
$\begin{split}\|\frac{1}{n_{l}}\mathbf{1}_{\mathcal{C}_{l}}^{T}\tilde{Z}_{[n_{C}]}\|_{2}^{2}&\leq\sum_{k\in[K]}\frac{C^{*}\log(n)}{n_{\mathcal{C}_{l}}N}=C^{*}\frac{K}{n_{\mathcal{C}_{l}}N}\log(n).\end{split}$
@@
Appendix D Synthetic Experiments
@@
    Generate matrices $W$ and $A$ For each cluster, we generate a topic mixture weight $\mathbf{\alpha}\sim\mathrm{Dirichlet}(\mathbf{u})$ where $u_{k}\sim\mathrm{Unif}(0.1,0.5)$ ($k\in[K]$). We order $\mathbf{\alpha}$ so that the biggest element of $\mathbf{\alpha}$ is assigned to the cluster’s dominant topic. We also add small Gaussian noise to $\alpha$ so that in the end, for each document in the cluster, $W_{i\cdot}=\mathbf{\alpha}+\mathbf{\epsilon}$, $\mathbf{\epsilon}_{k}\sim N(0,0.03)$. We sample $K$ rows of $W$ as anchor documents and set them as $\mathbf{e}_{k}$. The elements of $A$ are generated from $\mathrm{Unif}(0,1)$ and normalized so that each row of $A$ sums up to 1. Similarly to anchor documents, $K$ columns of $A$ are selected as anchor words and set to $\mathbf{e}_{k}$.
@@
We show the $\ell_{1}$ errors of estimated $\widehat{W}$ and $\widehat{A}$ under the same parameter settings as Section 3.4. We observe that the behavior of $\ell_{1}$ error is similar to $\ell_{2}$ error in Section 3.4. GpLSI has the lowest $\ell_{1}$ errors of $\widehat{W}$, followed by LDA and one-step graph-aligned denoising. We also observe that GpLSI has the lowest $\ell_{1}$ errors of $\widehat{A}$.
Figure 12: $\ell_{1}$ error for the estimator $\widehat{W}$ (defined as $\text{min}_{P\in\mathcal{P}}\frac{1}{n}\|\widehat{W}-WP\|_{F}$) for different combinations of document length $N$ and vocabulary size $p$. Here, $n=1000$ and $K=3$.
Figure 13: $\ell_{1}$ error for the estimator $\widehat{A}$ (defined as $\text{min}_{P\in\mathcal{P}}\frac{1}{p}\|\widehat{A}-PA\|_{F}$) for different combinations of document length $N$ and vocabulary size $p$. Here, $n=1000$ and $K=3$.
Figure 14: Computation time for different combinations of document length $N$ and vocabulary size $p$. Here, $n=1000$ and $K=3$.
@@
Appendix E Experiments
In this section, we provide supplementary plots for our analysis on the real datasets discussed in Section 4.
E.1 Estimated tumor-immune microenvironment topic weights in Section 4.1.1
@@ Appendix E Experiments
E.2 Kaplan-Meier curves of different topic groups in Section 4.1.1
@@
E.3 Topics by top common ingredients in Section 4.2
We illustrate each topic with the top ten ingredients with the highest estimated weights. Compared to anchor ingredients in Section 4.2, we observe that there are more overlapping ingredients among topics. While the top ten and anchor ingredients for each topic in GpLSI and LDA reflect similar styles, it is difficult to match anchor ingredients to the top ten ingredients in pLSI because the top ten ingredients are too similar across topics.

### v2

@@ E.3.2 Technical Lemmas
$\begin{split}&\mathbb{P}\left[\exists k\in[K],\exists l\in[n_{C}]:\quad\frac{1}{n_{\mathcal{C}_{l}}N}|\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{N}\sum_{j=1}^{p}\eta_{jm}|>\sqrt{C^{*}\frac{\log(n)}{n_{\mathcal{C}_{l}}N}}\right]\\ &\leq 2Kn_{C}\exp\{-{C^{*}\log(n)}\}\\ &=\exp\{\log(2)+\log(K)+\log(n_{C})-C^{*}\log(n)\}\\ &\leq\exp\{-(C^{*}-3)\log(n)\},\end{split}$
Thus with probability $1-o(n^{-1})$, for all $k\in[K]$ and all $l\in[n_{C}]$:
$\begin{split}\|\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}\tilde{Z}_{[n_{\mathcal{C}_{l}}]\cdot}\|_{2}^{2}&\leq\sum_{k\in[K]}\frac{C^{*}\log(n)}{n_{\mathcal{C}_{l}}N}=C^{*}\frac{K}{n_{\mathcal{C}_{l}}N}\log(n).\end{split}$
@@
Lemma 17 (Concentration of $\|e_{i}^{\top}Z\|_{2}$ and $\|e_{i}^{\top}ZV\|_{2}$).
Let Assumptions 1-5 hold. Let $\tilde{Z}=ZV$, with $V\in\mathbb{R}^{p\times r}$ a projection matrix: $V^{T}V=I_{r}$, with $r$ a term that does not grow with $n$ or $p$ and $r\leq p$. Then with probability at least $1-o(n^{-1})$:
$\begin{split}\max_{i\in[n]}\|e_{i}^{\top}Z\|_{2}&\leq C_{1}\sqrt{\frac{K\log(n)}{N}}\\ \max_{i\in[n]}\|e_{i}^{\top}ZV\|_{2}&\leq C_{2}\sqrt{\frac{r\log(n)}{N}}\\ \end{split}$ (84)
Proof.
We first note that
$\begin{split}\|e_{i}^{\top}Z\|_{2}^{2}&=\sum_{j=1}^{p}Z_{ij}^{2}\end{split}$
$\forall j\in[p],\qquad{Z}_{ij}$ $=\frac{1}{N}\sum_{m=1}^{N}(T_{im}(j)-\mathbb{E}[T_{im}(j)])$
$=\frac{1}{N}\sum_{m=1}^{N}\eta_{im}\quad\text{ with }\eta_{im}=T_{im}(j)-\mathbb{E}[T_{im}(j)]$
Thus ${Z}_{ij}$ is a sum of $N$ centered independent variables.
Fix $j\in[p]$. We have: $\text{Var}(\eta_{im})=M_{ij}^{2}-M_{ij}\leq M_{ij}\leq h_{j}$ and
$\sum_{m=1}^{N}\text{Var}(\eta_{im})\leq Nh_{j}.$
Moreover,for each $m$, $|\eta_{im}|\leq 1.$ Thus, by Lemma 10:
$\begin{split}\mathbb{P}[|\frac{1}{N}\sum_{m=1}^{N}\eta_{im}|>t]&\leq 2e^{-\frac{Nt^{2}/2}{1+t/3}}\end{split}$
Choosing $t=C^{*}\sqrt{\frac{h_{j}\log(n)}{N}}$:
$\begin{split}\mathbb{P}[|\frac{1}{N}\sum_{m=1}^{N}\sum_{j=1}^{p}\eta_{jm}|>t]&\leq 2e^{-(C^{*})^{2}\frac{\log(n)/2}{1+\frac{C^{*}}{3}\sqrt{\frac{\log(n)}{h_{j}N}}}}\end{split}$
Therefore, by Assumption 5, $N>c_{\min}\log(n)$, then, with probability at least $1-o(n^{-1})$, $|Z_{ij}|^{2}\leq C^{*}\frac{h_{j}\log(n)}{N}.$ By a simple union bound:
$\begin{split}&\mathbb{P}[\exists j:|Z_{ij}|^{2}>C\frac{h_{j}\log(n)}{N}]\leq pe^{-C^{*}\log(n)}\\ \implies&\mathbb{P}[\max_{i\in[n]}\|e_{i}^{\top}Z\|_{2}^{2}>C\frac{K\log(n)}{N}]\leq npe^{-C^{*}\log(n)}\leq e^{-C^{*}\log(n)+\log(p)+\log(n)}\leq e^{-(C^{*}-2)\log(n)}\end{split}$
since $\sum_{j=1}^{p}h_{j}=K$. Also, since we assume that $\max(p,r)\ll n$, the result follows.
Similarly, denote $\tilde{Z}=ZV$,
$\forall k\in[r],\qquad{\tilde{Z}}_{ik}$ $=\frac{1}{N}\sum_{j=1}^{p}\sum_{m=1}^{N}(T_{im}(j)V_{jk}-\mathbb{E}[T_{im}(j)V_{jk}])$
$=\frac{1}{N}\sum_{m=1}^{N}\sum_{j=1}^{p}\eta_{jm}\quad\text{ with}\eta_{jm}=V_{jk}T_{im}(j)-\mathbb{E}[V_{jk}T_{im}(j)]$
Note that: $\sum_{j=1}^{p}(T_{im}(j)V_{jk}-\mathbb{E}[T_{im}(j)V_{jk}])=V_{j_{0}K}-\sum_{j=1}^{p}M_{ij}V_{jk}$ with probability $M_{ij_{0}},j_{0}\in[p].$
Thus ${\tilde{Z}}_{ik}$ is a sum of $N$ centered independent variables.
Fix $k\in[r]$. We have: $\text{Var}(\sum_{j=1}^{p}\eta_{jm})=\sum_{j=1}^{p}V_{jk}^{2}-(\sum_{j=1}^{p}V_{jk}M_{ij})^{2},$ and since $\sum_{j=1}^{p}V_{jk}^{2}=1$:
$\sum_{m=1}^{N}\text{Var}(\sum_{j=1}^{p}\eta_{jm})\leq N.$
Moreover,for each $m$,
$|\sum_{j=1}^{p}\eta_{jm}|\leq\max_{j}|V_{jk}|+\sum_{j=1}^{p}M_{ij}|V_{jk}|\leq 2\max_{j}|V_{jk}|\leq 2.$
Thus, by Lemma 10:
$\begin{split}\mathbb{P}[|\frac{1}{N}\sum_{m=1}^{N}\sum_{j=1}^{p}\eta_{jm}|>t]&\leq 2e^{-\frac{Nt^{2}/2}{1+2t/3}}\end{split}$
Choosing $t=C^{*}\sqrt{\frac{\log(n)}{N}}$:
$\begin{split}\mathbb{P}[|\frac{1}{N}\sum_{m=1}^{N}\sum_{j=1}^{p}\eta_{jm}|>t]&\leq 2e^{-(C^{*})^{2}\frac{\log(n)/2}{1+\frac{2}{3}C^{*}\sqrt{\frac{\log(n)}{N}}}}\end{split}$
Therefore, by Assumption 5, $N>c_{\min}\log(n)$, then, with probability at least $1-o(n^{-1})$, $|\tilde{Z}_{kj}|^{2}\leq C^{*}\frac{\log(n)}{N}.$ By a simple union bound:
$\begin{split}&\mathbb{P}[\exists k:|\tilde{Z}_{ik}|^{2}>C\frac{\log(n)}{N}]\leq re^{-C^{*}\log(n)}\\ \implies&\mathbb{P}[\max_{i\in[n]}\|e_{i}^{\top}\tilde{Z}\|_{2}^{2}>C\frac{r\log(n)}{N}]\leq rne^{-C^{*}\log(n)}\leq e^{-C^{*}\log(n)+\log(r)+\log(n)}\leq e^{-(C^{*}-2)\log(n)}\end{split}$
Since we assume that $\max(p,r)\ll n$, the result follows. ∎
Appendix F Synthetic Experiments
@@
    Generate matrices $W$ and $A$ For each cluster, we generate a topic mixture weight $\mathbf{\alpha}\sim\mathrm{Dirichlet}(\mathbf{u})$ where $u_{k}\sim\mathrm{Unif}(0.1,0.5)$ ($k\in[K]$). We order $\mathbf{\alpha}$ so that the biggest element of $\mathbf{\alpha}$ is assigned to the cluster’s dominant topic. We also add small Gaussian noise to $\alpha$ so that in the end, for each document in the cluster, $W_{i\cdot}=\mathbf{\alpha}+\mathbf{\epsilon}_{i}$, $\mathbf{\epsilon}_{ik}\sim N(0,0.03)$. We sample $K$ rows of $W$ as anchor documents and set them as $\mathbf{e}_{k}$. The elements of $A$ are generated from $\mathrm{Unif}(0,1)$ and normalized so that each row of $A$ sums up to 1. Similarly to anchor documents, $K$ columns of $A$ are selected as anchor words and set to $\mathbf{e}_{k}$.
@@
Figure 10: $\ell_{1}$ error for the estimator $\widehat{W}$ (defined as $\text{min}_{P\in\mathcal{P}}\frac{1}{n}\|\widehat{W}-WP\|_{11}$) for different combinations of document length $N$ and vocabulary size $p$. Here, $n=1000$ and $K=3$.
Figure 11: $\ell_{1}$ error for the estimator $\widehat{A}$ (defined as $\text{min}_{P\in\mathcal{P}}\frac{1}{p}\|\widehat{A}-PA\|_{11}$) for different combinations of document length $N$ and vocabulary size $p$. Here, $n=1000$ and $K=3$.
@@
Next, we show the errors of estimated $\widehat{W}$ and $\widehat{A}$ under the same parameter settings as Section 3.4 of the main manuscript. From Figure 9-11, GpLSI achieves the lowest errors of $\widehat{W}$ and $\widehat{A}$ in all parameter settings, followed by LDA. For the estimation of $A$, as highlighted in Remark 4, our rates and procedure is not optimal compared with existing results (see in particular Ke & Wang (2017), which achieves similar results to ours in Figure 2 in the main manuscript). However, compared to the procedure proposed by Klopp et al. (2021), the estimation error is considerably improved.
Appendix G Real data
In this section, we provide supplementary plots for our analysis on the real datasets discussed in Section 4 of the main manuscript.
G.1 Estimated tumor-immune microenvironment topic weights
@@
G.2 Kaplan-Meier curves of Stanford Colorectal Cancer dataset
@@
G.3 Topics by top common ingredients in What’s Cooking dataset
@@
We illustrate each topic with the top ten ingredients with the highest estimated weights (Figures 19-21) as well as anchor ingredients for pLSI and LDA (Figures 17-18). Compared to anchor ingredients, we observe that there are more overlapping ingredients among topics. While the top ten and anchor ingredients for each topic in GpLSI and LDA reflect similar styles, it is difficult to match anchor ingredients to the top ten ingredients in pLSI because the top ten ingredients are too similar across topics.